OpenStudy (anonymous):

Simplify: 6y + 5z + 8y – 3z ?

OpenStudy (anonymous):

Surely you can do this?

OpenStudy (anonymous):

14y+2z

OpenStudy (anonymous):

I don't mean to be rude but this is very basic algebra.

OpenStudy (anonymous):

or 2(7y+z)

OpenStudy (kendricklamar2014):

Let's simplify step-by-step. 6y+5z+8y−3z =6y+5z+8y+−3z

OpenStudy (anonymous):

oh no i know what you mean i just had problems with math all my life

OpenStudy (kendricklamar2014):

Combine Like Terms: =6y+5z+8y+−3z =(6y+8y)+(5z+−3z)

OpenStudy (arabpride):

Come on you guys - there are people on OS that are younger than yourself that are still learning - be nice please :3

OpenStudy (kendricklamar2014):

Yes, @arabpride is right, be nice with other people.

OpenStudy (arabpride):

It's all a matter of combining like terms @quentavess :D do you know what that means?

OpenStudy (anonymous):

sort of yeah @arabpride

OpenStudy (anonymous):

It wasn't my intention to be rude, as I said in my second post. People post here for different reasons, I was trying to find out the reason to best help him/her. @quentavess, it might help to translate this question to a real life example. Let y be the number of apples and z be the number of oranges. In my basket, I add 6 apples, then 5 oranges, then 8 apples, then I take away 3 oranges. So we've put in 6+8=14 apples, and 5-3=2 oranges. Earlier we said y was the number of apples and we have 14 of them, so 14y, and then we said z was the number of oranges and we have 2 of them, so 2z. In total, we have 14y and 2z, so the answer is: 14y+2z

OpenStudy (arabpride):

6y + 5z + 8y – 3z <combining like terms> 6y + 5z + 8y – 3z -6y -6y 5z + 2y - 3z +3z +3z 8z + 2y <this is as far as you can simplify> <combining like terms just means to put alike variables together to help you simplify your equation> :D

OpenStudy (anonymous):

ooooohh i see thanks so much :) @arabpride @tom982

OpenStudy (arabpride):

You're Welcome :3

OpenStudy (anonymous):

My pleasure, @quentavess. Good luck with the rest of your algebra!